The smallest angle in a triangle is 1/2 the size of the largest. The middle angle is 25 less than the largest. Find all three angles.
Thank you

Answer by algebrahouse.com(1659)

x = the largest angle
0.5x = smallest angle {smallest angle is half the largest}
x - 25 {middle angle is 25 less than the largest}

x + 0.5x + x - 25 = 180 {angles of a triangle add up to 180°}
2.5x - 25 = 180 {combined like terms}
2.5x = 205 {added 25 to each side}
x = 82 {divided each side by 2.5}
0.5x = 41 {substituted 82, for x, into 0.5x}
x - 25 = 57 {substituted 82, for x, into x - 25}

82 is the largest angle
57 is the middle angle
41 is the smallest angle

For more help from me, visit: www.algebrahouse.com